# Huan Zhang

### v1

@@ 4 An Inertial Accelerated Primal-dual Algorithm
$\begin{cases}\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\Bigg{(}\nabla f(x(t))+\partial g(x(t))+A^{*}\left(\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t)\right)+\rho A^{*}\left(Ax(t)-b\right)\Bigg{)}\ni 0,\\ \ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\beta(t)\left(A\left(x(t)+\frac{t}{\alpha-1}\dot{x}(t)\right)-b\right)=0.\end{cases}$
@@
$\begin{cases}u(t):=x(t)+\frac{t}{\alpha-1}\dot{x}(t),\\ v(t):=\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t).\end{cases}$
@@
$\begin{cases}u(t)=x(t)+\frac{t}{\alpha-1}\dot{x}(t),\\ \dot{u}(t)\in-\frac{t}{\alpha-1}\beta(t)\left(\nabla f(x(t))+\partial g(x(t))\right)-\frac{t}{\alpha-1}\beta(t)A^{*}v(t)-\frac{t}{\alpha-1}\rho\beta(t)A^{*}(Ax(t)-b),\\ v(t)=\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t),\\ \dot{v}(t)=\frac{t}{\alpha-1}\beta(t)\left(Au(t)-b\right).\end{cases}$
@@
$\begin{cases}u_{k+1}=x_{k+1}+\frac{k}{\alpha-1}(x_{k+1}-x_{k}),\\ \widetilde{u}_{k+1}-u_{k}\in-\frac{k}{\alpha-1}\beta_{k}\left(\nabla f(\widetilde{x}_{k+1})+\partial g(x_{k+1})\right)-\frac{k}{\alpha-1}\beta_{k}A^{*}v_{k+1}-\frac{k}{\alpha-1}\rho\beta_{k}A^{*}(Ax_{k+1}-b),\end{cases}$
@@
$\begin{cases}u_{k+1}=x_{k+1}+\frac{k}{\alpha-1}(x_{k+1}-x_{k}),\\ u_{k+1}-u_{k}\in-\frac{k+\alpha-1}{\alpha-1}\beta_{k}\left(\nabla f(\widetilde{x}_{k+1})+\partial g(x_{k+1})\right)-\frac{k+\alpha-1}{\alpha-1}\beta_{k}A^{*}v_{k+1}-\frac{k+\alpha-1}{\alpha-1}\rho\beta_{k}A^{*}(Ax_{k+1}-b).\end{cases}$
@@
$\begin{cases}v_{k+1}=\lambda_{k+1}+\frac{\sqrt{\sigma_{k}}k}{\alpha-1}\frac{\lambda_{k+1}-\lambda_{k}}{\sqrt{\sigma_{k}}},\\ \frac{v_{k+1}-v_{k}}{\sqrt{\sigma_{k}}}=\frac{\sqrt{\sigma_{k}}k}{\alpha-1}\beta_{k}(Au_{k+1}-b).\end{cases}$ (19)
@@
$\begin{cases}u_{k+1}=x_{k+1}+(t_{k+1}-1)(x_{k+1}-x_{k}),\\ u_{k+1}-u_{k}\in-t_{k+1}\beta_{k}\left(\nabla f(\widetilde{x}_{k+1})+\partial g(x_{k+1})\right)-t_{k+1}\beta_{k}A^{*}v_{k+1}-\rho t_{k+1}\beta_{k}A^{*}(Ax_{k+1}-b),\\ v_{k+1}=\lambda_{k+1}+(t_{k+1}-1)(\lambda_{k+1}-\lambda_{k}),\\ \mu_{k}=\lambda_{k}+\frac{t_{k}-1}{t_{k+1}}(\lambda_{k}-\lambda_{k-1}),\\ \lambda_{k+1}=\mu_{k}+\sigma\beta_{k}(Au_{k+1}-b).\end{cases}$
@@
$v_{k+1}$ $=t_{k+1}\lambda_{k+1}-(t_{k+1}-1)\lambda_{k}$
$=t_{k+1}\mu_{k}+\sigma t_{k+1}\beta_{k}(Au_{k+1}-b)-(t_{k+1}-1)\lambda_{k}$
$=t_{k+1}\mu_{k}+\sigma t_{k+1}\beta_{k}\left(t_{k+1}Ax_{k+1}-(t_{k+1}-1)Ax_{k}-b\right)-(t_{k+1}-1)\lambda_{k}$
$=t_{k+1}\mu_{k}-(t_{k+1}-1)\lambda_{k}+\sigma\beta_{k}t_{k+1}^{2}\left(Ax_{k+1}-\frac{1}{t_{k+1}}\left((t_{k+1}-1)Ax_{k}+b\right)\right).$
@@
$\begin{aligned} &\frac{1}{2}\lVert u_{k+1}-x^{*}\rVert^{2}-\frac{1}{2}\lVert u_{k}-x^{*}\rVert^{2}\\ &=\left\langle u_{k+1}-u_{k},u_{k+1}-x^{*}\right\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}\\ &=\left\langle t_{k+1}(x_{k+1}-x_{k})-(t_{k}-1)(x_{k}-x_{k-1}),\right.\\ &~~~~\left.x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}\\ &=-t_{k+1}\beta_{k}\bigg{\langle}\nabla f(\bar{x}_{k})+\eta_{k+1}+A^{*}v_{k+1}+\rho A^{*}(Ax_{k+1}-b),\\ &~~~~x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\bigg{\rangle}-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}\\ &=-t_{k+1}\beta_{k}\left\langle\nabla f(\bar{x}_{k}),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle\\ &~~~-t_{k+1}\beta_{k}\left\langle A^{*}v_{k+1}+\eta_{k+1},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle\\ &~~~-t_{k+1}\beta_{k}\left\langle\rho A^{*}(Ax_{k+1}-b),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}\\ &=-t_{k+1}\beta_{k}\left\langle\nabla f(\bar{x}_{k}),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle\\ &~~~-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x^{*}\right\rangle\\ &~~~-t_{k+1}(t_{k+1}-1)\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x_{k}\right\rangle\\ &~~~-\rho t_{k+1}\beta_{k}\left\langle A^{*}(Ax_{k+1}-b),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2},\end{aligned}$
@@
$\begin{aligned} &-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle\\ &=-t_{k+1}\beta_{k}\left\langle A^{*}v_{k+1}+\eta_{k+1}-\eta_{k+1}-A^{*}\lambda^{*},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle\\ &=-t_{k+1}\beta_{k}\left\langle A^{*}v_{k+1}+\eta_{k+1},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle\\ &~~~~+t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x^{*}\right\rangle+t_{k+1}(t_{k+1}-1)\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x_{k}\right\rangle.\end{aligned}$
@@
$~~~~+\frac{L_{f}\beta_{k}t_{k+1}^{2}}{2}\left\|x_{k+1}-\bar{x}_{k}\right\|^{2}$
@@
$~~~~+\frac{L_{f}\beta_{k}}{2}\left\|u_{k+1}-u_{k}\right\|^{2}.$
@@
$\begin{aligned} &-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x^{*}\right\rangle-t_{k+1}(t_{k+1}-1)\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x_{k}\right\rangle\\ &\leq-t_{k+1}\beta_{k}\left(g(x_{k+1})-g(x^{*})\right)-t_{k+1}(t_{k+1}-1)\beta_{k}\left(g(x_{k+1})-g(x_{k})\right)\\ &~~-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle.\end{aligned}$
@@
$~~~~-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-Ax_{k}\right\|^{2}$
@@
$~~~~-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}.$
@@
$\begin{aligned} &E_{1}(k+1)-E_{1}(k)\\ &=\frac{1}{2}\lVert u_{k+1}-x^{*}\rVert^{2}-\frac{1}{2}\lVert u_{k}-x^{*}\rVert^{2}\\ &\leq-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle-t_{k+1}\beta_{k}\left(f(x_{k+1})+g(x_{k+1})-f(x^{*})-g(x^{*})\right)\\ &~~~~-t_{k+1}(t_{k+1}-1)\beta_{k}\left(f(x_{k+1})+g(x_{k+1})-f(x_{k})-g(x_{k})\right)\\ &~~~~-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle-\rho t_{k+1}\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}\\ &~~~~+\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k}-b\right\|^{2}-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}\\ &~~~~+\frac{L_{f}\beta_{k}-1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}\\ &\leq-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle-t_{k+1}\beta_{k}\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~~~-t_{k+1}(t_{k+1}-1)\beta_{k}\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x_{k},\lambda^{*})\right)+\frac{L_{f}\beta_{k}-1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}.\end{aligned}$
@@
$E_{2}(k+1)-E_{2}(k)$ $=\frac{1}{2\sigma}\lVert v_{k+1}-\lambda^{*}\rVert^{2}-\frac{1}{2\sigma}\lVert v_{k}-\lambda^{*}\rVert^{2}$
@@
$E(k+1)-E(k)$ (37)
$=E_{0}(k+1)-E_{0}(k)+E_{1}(k+1)-E_{1}(k)+E_{2}(k+1)-E_{2}(k)$
$\leq\left(t_{k+2}(t_{k+2}-1)\beta_{k+1}-t_{k+1}^{2}\beta_{k}\right)\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$
$~~+\frac{L_{f}\beta_{k}-1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}-\frac{1}{2\sigma}\lVert v_{k+1}-v_{k}\rVert^{2}.$
@@
$\begin{aligned} &\sum_{k=1}^{N}(t_{k+1}^{2}\beta_{k}-t_{k+2}(t_{k+2}-1)\beta_{k+1})\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &~~+\sum_{k=1}^{N}\frac{1-L_{f}\beta_{k}}{2}\left\|u_{k+1}-u_{k}\right\|^{2}+\frac{1}{2\sigma}\sum_{k=1}^{N}\lVert v_{k+1}-v_{k}\rVert^{2}\leq E(1)-E(N+1)\leq E(1).\end{aligned}$
@@
Let $\{(x_{k},\lambda_{k})\}_{k\geq 0}$ be the sequence generated by $\mathrm{IAPDA}$ and let $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$. Then, for each $k\geq 1$,
$\mathcal{L}_{\rho}(x_{k},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\leq\frac{E(1)}{t_{k+1}(t_{k+1}-1)\beta_{k}},~~\|Ax_{k}-b\|\leq\frac{\beta_{0}t_{1}^{2}\|Ax_{1}-b\|+2C}{\beta_{k}t_{k+1}(t_{k+1}-1)},$
@@
$~~~~+\frac{\rho}{2}\left(\frac{\beta_{0}t_{1}^{2}\|Ax_{1}-b\|+2C}{\beta_{k}t_{k+1}(t_{k+1}-1)}\right)^{2},$
@@
$~~~~+\frac{\rho}{2}\left(\frac{\beta_{0}t_{1}^{2}\|Ax_{1}-b\|+2C}{\beta_{k}t_{k+1}(t_{k+1}-1)}\right)^{2}.$

### v2

@@ 4 An Inertial Accelerated Primal-dual Algorithm
$\left\{\begin{split}&\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\Bigg(\nabla f(x(t))+\partial g(x(t))\\ &~~~~~~~~~~~~~~+A^{*}\left(\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t)\right)+\rho A^{*}\left(Ax(t)-b\right)\Bigg)\ni 0,\\ &\ddot{\lambda}(t)+\frac{\alpha}{t}\dot{\lambda}(t)-\beta(t)\left(A\left(x(t)+\frac{t}{\alpha-1}\dot{x}(t)\right)-b\right)=0.\end{split}\right$. (12)
@@
$\left\{\begin{split}&u(t):=x(t)+\frac{t}{\alpha-1}\dot{x}(t),\\ &v(t):=\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t).\end{split}\right$.
@@
$\left\{\begin{split}&u(t)=x(t)+\frac{t}{\alpha-1}\dot{x}(t),\\ &\dot{u}(t)\in-\frac{t}{\alpha-1}\beta(t)\left(\nabla f(x(t))+\partial g(x(t))\right)-\frac{t}{\alpha-1}\beta(t)A^{*}v(t)\\ &~~~~~~~~-\frac{t}{\alpha-1}\rho\beta(t)A^{*}(Ax(t)-b),\\ &v(t)=\lambda(t)+\frac{t}{\alpha-1}\dot{\lambda}(t),\\ &\dot{v}(t)=\frac{t}{\alpha-1}\beta(t)\left(Au(t)-b\right).\end{split}\right$. (13)
@@
$\left\{\begin{split}&u_{k+1}=x_{k+1}+\frac{k}{\alpha-1}(x_{k+1}-x_{k}),\\ &\widetilde{u}_{k+1}-u_{k}\in-\frac{k}{\alpha-1}\beta_{k}\left(\nabla f(\widetilde{x}_{k+1})+\partial g(x_{k+1})\right)-\frac{k}{\alpha-1}\beta_{k}A^{*}v_{k+1}\\ &~~~~~~~~~~~~~~~~~~-\frac{k}{\alpha-1}\rho\beta_{k}A^{*}(Ax_{k+1}-b),\end{split}\right$. (14)
@@
$\left\{\begin{split}&u_{k+1}=x_{k+1}+\frac{k}{\alpha-1}(x_{k+1}-x_{k}),\\ &u_{k+1}-u_{k}\in-\frac{k+\alpha-1}{\alpha-1}\beta_{k}\left(\nabla f(\widetilde{x}_{k+1})+\partial g(x_{k+1})\right)-\frac{k+\alpha-1}{\alpha-1}\beta_{k}A^{*}v_{k+1}\\ &~~~~~~~~~~~~~~~~-\frac{k+\alpha-1}{\alpha-1}\rho\beta_{k}A^{*}(Ax_{k+1}-b).\end{split}\right$.
@@
$\left\{\begin{split}v_{k+1}=\lambda_{k+1}+\frac{\sqrt{\sigma_{k}}k}{\alpha-1}\frac{\lambda_{k+1}-\lambda_{k}}{\sqrt{\sigma_{k}}},\\ \frac{v_{k+1}-v_{k}}{\sqrt{\sigma_{k}}}=\frac{\sqrt{\sigma_{k}}k}{\alpha-1}\beta_{k}(Au_{k+1}-b).\end{split}\right$. (15)
@@
$\left\{\begin{split}&u_{k+1}=x_{k+1}+(t_{k+1}-1)(x_{k+1}-x_{k}),\\ &u_{k+1}-u_{k}\in-t_{k+1}\beta_{k}\left(\nabla f(\widetilde{x}_{k+1})+\partial g(x_{k+1})\right)-t_{k+1}\beta_{k}A^{*}v_{k+1}\\ &~~~~~~~~~~~~~~~~-\rho t_{k+1}\beta_{k}A^{*}(Ax_{k+1}-b),\\ &v_{k+1}=\lambda_{k+1}+(t_{k+1}-1)(\lambda_{k+1}-\lambda_{k}),\\ &\mu_{k}=\lambda_{k}+\frac{t_{k}-1}{t_{k+1}}(\lambda_{k}-\lambda_{k-1}),\\ &\lambda_{k+1}=\mu_{k}+\sigma\beta_{k}(Au_{k+1}-b).\end{split}\right$. (16)
@@
$\begin{split}v_{k+1}&=t_{k+1}\lambda_{k+1}-(t_{k+1}-1)\lambda_{k}\\ &=t_{k+1}\mu_{k}+\sigma t_{k+1}\beta_{k}(Au_{k+1}-b)-(t_{k+1}-1)\lambda_{k}\\ &=t_{k+1}\mu_{k}+\sigma t_{k+1}\beta_{k}\left(t_{k+1}Ax_{k+1}-(t_{k+1}-1)Ax_{k}-b\right)-(t_{k+1}-1)\lambda_{k}\\ &=t_{k+1}\mu_{k}-(t_{k+1}-1)\lambda_{k}+\sigma\beta_{k}t_{k+1}^{2}\left(Ax_{k+1}-\frac{1}{t_{k+1}}\left((t_{k+1}-1)Ax_{k}+b\right)\right).\end{split}$
@@
Remark 4.2.
It is worth noting that, in practice, it is difficult to obtain an exact solution to the subproblem ($\mathbf{Step~2}$ in IAPDA). Accordingly, several algorithms have been developed to compute inexact solutions for such subproblems in practical problem-solving scenarios. Specifically, one can employ some classical splitting methods such as the proximal algorithm and the accelerated scheme FISTA b2020 ; beck2009 , as well as inexact versions derived from continuous dynamics to discrete algorithms He2025 ; he2026 .
@@
$\frac{1}{2}\lVert u_{k+1}-x^{*}\rVert^{2}-\frac{1}{2}\lVert u_{k}-x^{*}\rVert^{2}$ (30)
$=$ $\left\langle u_{k+1}-u_{k},u_{k+1}-x^{*}\right\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}$
$=$ $\left\langle t_{k+1}(x_{k+1}-x_{k})-(t_{k}-1)(x_{k}-x_{k-1}),\right$.
$\left.x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}$
$=$ $-t_{k+1}\beta_{k}\bigg\langle\nabla f(\bar{x}_{k})+\eta_{k+1}+A^{*}v_{k+1}+\rho A^{*}(Ax_{k+1}-b),$
$x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\bigg\rangle-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}$
$=$ $-t_{k+1}\beta_{k}\left\langle\nabla f(\bar{x}_{k}),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$-t_{k+1}\beta_{k}\left\langle A^{*}v_{k+1}+\eta_{k+1},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$-t_{k+1}\beta_{k}\left\langle\rho A^{*}(Ax_{k+1}-b),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}$
$=$ $-t_{k+1}\beta_{k}\left\langle\nabla f(\bar{x}_{k}),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x^{*}\right\rangle$
$-t_{k+1}(t_{k+1}-1)\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x_{k}\right\rangle$
$-\rho t_{k+1}\beta_{k}\left\langle A^{*}(Ax_{k+1}-b),x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$-\frac{1}{2}\left\|u_{k+1}-u_{k}\right\|^{2},$
@@
$-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle$
$=$ $-t_{k+1}\beta_{k}\left\langle A^{*}v_{k+1}+\eta_{k+1}-\eta_{k+1}-A^{*}\lambda^{*},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$=$ $-t_{k+1}\beta_{k}\left\langle A^{*}v_{k+1}+\eta_{k+1},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle$
$+t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x^{*}\right\rangle+t_{k+1}(t_{k+1}-1)\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x_{k}\right\rangle.$
@@
$+\frac{L_{f}\beta_{k}t_{k+1}^{2}}{2}\left\|x_{k+1}-\bar{x}_{k}\right\|^{2}$
@@
$+\frac{L_{f}\beta_{k}}{2}\left\|u_{k+1}-u_{k}\right\|^{2}.$
@@
$-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x^{*}\right\rangle-t_{k+1}(t_{k+1}-1)\beta_{k}\left\langle A^{*}\lambda^{*}+\eta_{k+1},x_{k+1}-x_{k}\right\rangle$ (32)
$\leq$ $-t_{k+1}\beta_{k}\left(g(x_{k+1})-g(x^{*})\right)-t_{k+1}(t_{k+1}-1)\beta_{k}\left(g(x_{k+1})-g(x_{k})\right)$
$-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle.$
@@
$-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-Ax_{k}\right\|^{2}$
@@
$-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}.$
@@
$E_{1}(k+1)-E_{1}(k)$
$=$ $\frac{1}{2}\lVert u_{k+1}-x^{*}\rVert^{2}-\frac{1}{2}\lVert u_{k}-x^{*}\rVert^{2}$
$\leq$ $-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle-t_{k+1}\beta_{k}\Big(f(x_{k+1})+g(x_{k+1})-f(x^{*})-g(x^{*})\Big)$
$-t_{k+1}(t_{k+1}-1)\beta_{k}\Big(f(x_{k+1})+g(x_{k+1})-f(x_{k})-g(x_{k})\Big)$
$-t_{k+1}\beta_{k}\left\langle A^{*}\lambda^{*},x_{k+1}-x^{*}+(t_{k+1}-1)(x_{k+1}-x_{k})\right\rangle-\rho t_{k+1}\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}$
$+\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k}-b\right\|^{2}-\frac{\rho}{2}t_{k+1}(t_{k+1}-1)\beta_{k}\left\|Ax_{k+1}-b\right\|^{2}$
$+\frac{L_{f}\beta_{k}-1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}$
$\leq$ $-t_{k+1}\beta_{k}\left\langle A^{*}(v_{k+1}-\lambda^{*}),u_{k+1}-x^{*}\right\rangle-t_{k+1}\beta_{k}\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$
$-t_{k+1}(t_{k+1}-1)\beta_{k}\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x_{k},\lambda^{*})\right)+\frac{L_{f}\beta_{k}-1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}.$
@@
$E_{2}(k+1)-E_{2}(k)$
$=$ $\frac{1}{2\sigma}\lVert v_{k+1}-\lambda^{*}\rVert^{2}-\frac{1}{2\sigma}\lVert v_{k}-\lambda^{*}\rVert^{2}$
@@
$\begin{split}&E(k+1)-E(k)\\ =&E_{0}(k+1)-E_{0}(k)+E_{1}(k+1)-E_{1}(k)+E_{2}(k+1)-E_{2}(k)\\ \leq&\left(t_{k+2}(t_{k+2}-1)\beta_{k+1}-t_{k+1}^{2}\beta_{k}\right)\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)\\ &+\frac{L_{f}\beta_{k}-1}{2}\left\|u_{k+1}-u_{k}\right\|^{2}-\frac{1}{2\sigma}\lVert v_{k+1}-v_{k}\rVert^{2}.\end{split}$ (33)
@@
$\sum_{k=1}^{N}(t_{k+1}^{2}\beta_{k}-t_{k+2}(t_{k+2}-1)\beta_{k+1})\left(\mathcal{L}_{\rho}(x_{k+1},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\right)$
$~~+\sum_{k=1}^{N}\frac{1-L_{f}\beta_{k}}{2}\left\|u_{k+1}-u_{k}\right\|^{2}+\frac{1}{2\sigma}\sum_{k=1}^{N}\lVert v_{k+1}-v_{k}\rVert^{2}$
$\leq E(1)-E(N+1)\leq E(1).$
@@
Let $\{(x_{k},\lambda_{k})\}_{k\geq 0}$ be the sequence generated by $\mathrm{IAPDA}$ and let $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$. Then, for each $k\geq 1$, it holds that
$\mathcal{L}_{\rho}(x_{k},\lambda^{*})-\mathcal{L}_{\rho}(x^{*},\lambda^{*})\leq\frac{E(1)}{t_{k+1}(t_{k+1}-1)\beta_{k}},$
$\|Ax_{k}-b\|\leq\frac{\beta_{0}t_{1}^{2}\|Ax_{1}-b\|+2C}{\beta_{k}t_{k+1}(t_{k+1}-1)},$
@@
$+\frac{\rho}{2}\left(\frac{\beta_{0}t_{1}^{2}\|Ax_{1}-b\|+2C}{\beta_{k}t_{k+1}(t_{k+1}-1)}\right)^{2},$
@@
$+\frac{\rho}{2}\left(\frac{\beta_{0}t_{1}^{2}\|Ax_{1}-b\|+2C}{\beta_{k}t_{k+1}(t_{k+1}-1)}\right)^{2}.$